DETECTION OF TWO HIGHLY-STABLE SILICON NITRIDES BY CHIRPED-PULSE ROTATIONAL SPECTROSCOPY: HSiNSi and SiH_{3}NSi

\underline{MICHAEL C. McCARTHY}, KYLE N. CRABTREE, OSCAR MARTINEZ, JR., Harvard-Smithsonian Center for Astrophysics, 60 Garden Street, Cambridge, MA 02138, and School of Engineering and Applied Sciences, Harvard University, 29 Oxford Street, Cambridge, MA 02138.

By means of chirped-pulse and cavity Fourier transform microwave spectroscopy of a supersonic molecular beam, the rotational spectra of two new silicon nitrides, HSiNSi and SiH_{3}NSi, have been detected in a discharge through dilute gas mixtures of either molecular nitrogen or ammonia and silane. Both molecules appear to be highly stable: they are by far the most readily observed silicon-nitrogen containing molecules in the 6-20 GHz frequency range, even though neither has apparently been the subject of prior experimental or theoretical studies. Density-functional calculations performed here confirm that both structures are deep minima on their respective potential energy surfaces, and each possesses a sizable dipole moment. Owing to their high abundance, in excess of 10^{12} molecules/gas pulse, extensive isotopic spectroscopy has been undertaken for both. Detection of other silicon-nitride clusters in the transition region from atomic constituents to Si_{3}N_{4} by similar means may be promising.